What is entropy ? Calculate the change in entropy of an ideal gas for an isothermal expansion.
1
Expert's answer
2019-05-20T14:02:47-0400

entropy is an extensive property of a thermodynamic system: the entropy S is the natural logarithm of the number of microstates, multiplied by the Boltzmann constant kB.

the change in entropy of an ideal gas for an isothermal expansion:


"S = \int \frac{\delta Q}{T} = \frac{1}{T} \int \delta Q = \frac{Q}{T}"

where T is temperature of an idela gas, Q is the heat absorbed during the isothermal expansion